8/x +24, given that x=-5
8/-5 + 24 = 22.4
the answer is 22.4 in the back of the book. How do I deal with the fraction with a negative bottom?
Thanks for your help.

Answer by rfer(16322)

8/-5+24=
-1.6+24=22.4